OpenStudy (anonymous):

Irma needs to rent a car for a business trip from Phoenix to San Jose (720 miles one way). She will drive round trip and will be gone a total of 8 days. The car rental company charges $ 65.00 per day and allows for 1,600 miles for Irmas's trip. There will be an additional charge of $ 0.24 per mile over the 1,600 miles allowed. How many total miles can she drive without exceeding her budget of $ 600.00?

OpenStudy (snowsurf):

Total cost is going to come from the rental fees and if she goes over the 1600 miles. Rental fee is $ 65/day. She rents the car for 8 days so 65 * 8 = 520 dollars. That is just the rental fees. We want to know how many miles should could drive without exceeding her budget. We got 520 dollars from rental fees. 600 dollars is her total budget for everything. So we take 600 - 520 = 80. So she would have 80 dollars she could spend for additional miles over 1600 miles. So 0.24x=80 and we want to find miles. x =80/0.24, x =333.3 miles. So under her budget she can add additional 333.3 miles to her rental car. So that will give her a total miles she can use of 1600 + 333.3 = 1933.3 miles.

OpenStudy (anonymous):

Thanks for answering :)